An object is short from the ground at 125m/s at an angle of 30 degree above the horizantal.How far away does the Land.
1
Expert's answer
2019-09-12T10:39:39-0400

"v_0 = 125\frac{\text{m}}{\text{s}}; \alpha = 30\degree; \\
d = \frac{v_0^2\sin{2\alpha}}{g} = \frac{(125\frac{\text{m}}{\text{s}})^2\sin{(2\cdot30\degree)}}{9.81\frac{\text{m}}{\text{s}^2}} \approx 1379\text{ m}."

For derivations, you may visit:

https://en.wikipedia.org/wiki/Range_of_a_projectile